OpenStudy (anonymous):

49/56 = 21/F how do i solve this please help

OpenStudy (a1234):

56*21 = 1176 1176/49 = ?

OpenStudy (calculusxy):

49/56 can be simplified to 7/8 You have the cross mutliply to solve for f since this is proportions. 7. 21 --- = ----- 8. F Mutliply 21*8=168 168=7F Divide 168 by 7 and you should get the value of f.

OpenStudy (calculusxy):

Do you understand?

OpenStudy (anonymous):

yes thank you very much!!

OpenStudy (calculusxy):

No problem :)